We are using a new version of the 'Alopeke control software. Be aware that the FITS headers have changed slightly.
After the short thermal cycle on Dec 18, 2019 to replace the coldhead's displacer, the charge smearing in GMOS-S CCD1 has re-appeared as shown in the screenshot of the N&S dark from April 2019. Any programs using the N&S mode are affected and the programs set up with IFU-2 or IFU-R observations are possibly affected by this as well. PIs of those programs are contacted to assess the impact on their data quality and the programs are set on hold. A controlled thermal cycle is scheduled to resolve the issue at the end of January.
The GSAOI web pages have been updated to reflect the recent changes introduced by the replacement of the guiding unit in Canopus. The new guiding unit called Natural Guide Star Next Generation Sensor (NGSNGS or NGS2) has been installed in July 2019. The on-sky characterization of NGS2 have been taken in place in October and December 2019. Results from commissioning show an improvement in sensitivity of about ~2.5 magnitudes, compared to the previous NGS guiding unit. In addition, there is a large improvement in acquisition and offsetting.
To highlight relevant issues that are still relevant I have moved old/historical information to the end.
Update June 26th, 2019
As a friendly reminder proposers should be aware that:
NIFS has been available since semester 2006B for classical and queue scheduled observing at Gemini-North.
January 2013: On Instrument Wave Front Sensor (OIWFS) is unavailable due to a faulty sensor.
October 2012: New 'LGS+P1' guide mode is now operational, allowing 'seeing improver' AO correction with essentially 100% sky coverage.
2018 Nov 26 - Dec 14
GNIRS was removed from the telescope to work on filter wheel #2 which had been showing increasing instances of not going into position reliably. The filter wheel datum switches were replaced on both filter mechanisms, a temporary light baffle was removed from wheel #1, and dust was removed from the filters using CO2 snow.
2018 Jun 20-25
Re-seated the detector controller boards, replaced the sequencer board, and replaced a voltage regulator. The read noise levels are now back to normal.
2018 Jun 5
Update November 6th, 2019
F2 is back on sky, fully operational for imaging and long-slit spectroscopy with the OIWFS.
Update September 26th, 2019
Recovery of the cooling was unsuccessful and the cold-heads in F2 must be replaced. This means that the instrument will be unavailable for approximately 4 weeks.
Update September 12th, 2019
The F2 detector has been warming up and has reached the limit for operation. We are working to recover the cooling, but until this is fixed F2 is not available for observations.
Sep 16, 2019
April 2, 2019
2018 Oct 17